OpenStudy (anonymous):

What is the equation of the line that has a slope of -1 and goes through point (0, 5)? Answer y = -x + 5 y = -2x + 5 y = -x y = 2x + 5

OpenStudy (anonymous):

@jennychan12 @abcabc

OpenStudy (jennychan12):

i would use pint slope form Y - 5 = -(x) Y = -x+5 A <--ANSWER

OpenStudy (anonymous):

Hey, Mr. Kool. I love multiple choice questions because you can do process of elimination. We know the equation needs a slope of -1. So only the first and third choice qualify. Then, just mentally substitute (0,5) in the second and fourth choice to see which ends up giving a true statement. Yup, it's the first of the four choices! Done! Hope that helped.

OpenStudy (jennychan12):

point-slope form (use when given a point and a slope) \[(Y-Y _{1}) = m(X-X _{1})\] slope intercept form y = mx+b